a an ice chest contains six cans of apple juice, eight cans of grape juice, four cans of orange juice, ans 2 cans of mango juice. suppose that you reach into the container and randomly select three cans in succession without replacement. find the probability of selecting three cans of apple juice.
Answer by jorel1380(3719)

6+8+4+2=20 total cans
6/20x5/19x4/18=120/6840=3/171..